# End-to-End Dexterous Arm-Hand VLA Policies via Shared Autonomy: VR Teleoperation Augmented by Autonomous Hand VLA Policy for Efficient Data Collection

### v1

@@ 8.3.2 Learning Industrial Assembly Task - Data Collection through Motion Planning
8.4 Summary and Discussion
The collective evidence from all experiments – spanning pick-and-place tasks with shared autonomy (main text), long-horizon tasks with teleoperation, and industrial assembly with motion planning – consistently demonstrates the effectiveness of our corrective framework across task types and data collection approaches. Each study followed the same fundamental methodology: initial policy training, failure identification, corrective data collection, and policy re-training/refinement, yet employed different failure-correction methods and data sources tailored to task requirements.

### v2

@@ 8.3.2 Learning Industrial Assembly Task - Data Collection through Motion Planning
8.4 Discussion and Summary
The collective evidence from all experiments – spanning pick-and-place tasks with shared autonomy (main text), long-horizon tasks with teleoperation, and industrial assembly with motion planning – consistently demonstrates the effectiveness of our corrective framework across task types and data collection approaches. Each study followed the same fundamental principle: initial policy training, failure identification, corrective data collection, and policy re-training/refinement, while employing different failure-correction approaches and data sources tailored to the task-specific requirements. This iterative approach, despite relying on manual intervention and human involvement, represents a preliminary implementation of a closed-loop data flywheel that continuously retrains our VLA models with real-world interactions and the correspondingly generated real-robot data.